undeadknight26 (undeadknight26):

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 7 over 8 power, all over 2 to the 1 over 4 power the eighth root of 2 to the fifth power the fifth root of 2 to the eighth power the square root of 2 to the 5 over 8 power the fourth root of 2 to the sixth power

undeadknight26 (undeadknight26):

@SolomonZelman @Hero please help!

undeadknight26 (undeadknight26):

@ash2326 @bigeyes420 @Compassionate please help!

OpenStudy (anonymous):

how did u do that tagg me in it

OpenStudy (campbell_st):

the index law for division of the same base is subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] apply this to your question

undeadknight26 (undeadknight26):

the at sign and then a name

undeadknight26 (undeadknight26):

But first i would need common denominators right>

undeadknight26 (undeadknight26):

so its the second one?

OpenStudy (campbell_st):

nope , you have \[2^{\frac{7}{8} - \frac{1}{4}} = \]

undeadknight26 (undeadknight26):

5 over 8?

OpenStudy (campbell_st):

then you need to know about indexes for radicals \[\sqrt[a]{x^b}=x^{\frac{b}{a}}\] now apply this to you part solution \[2^{\frac{5}{8}}\]

undeadknight26 (undeadknight26):

8/2^5?

OpenStudy (campbell_st):

I think that makes sense... which answer choice would it be..?

undeadknight26 (undeadknight26):

The first one?

OpenStudy (campbell_st):

yep thats it

undeadknight26 (undeadknight26):

Can you help me with 4 more please?

undeadknight26 (undeadknight26):

or is that 2 much?

OpenStudy (campbell_st):

sorry I'm about to head off and do some jobs...

undeadknight26 (undeadknight26):

Ok thanks! Go help the world!

OpenStudy (anonymous):

Did that end up being the right answer? @undeadknight26

OpenStudy (anonymous):

was this correct